A mass m=2.50kg is being accelerated upwards by tension in a string. The magnitude of the acceleration a=4.00m/s2. Calculate the magnitude of the tension in the string


1
Expert's answer
2021-07-12T11:32:02-0400

Gives

Mass(m)=2.50kg

acceleration (a)=4.00m/sec2

accelerated upwards by tension in a string. 

"T-mg=ma"


"T=m(g+a)=2.50\times(9.8+4)=34.5N"